what is the max kinetic energy of a longitudnal wave in terms of density and frequency??

what is the  max kinetic energy of a longitudnal wave in terms of density and frequency??
 

Grade:11

1 Answers

Norman Manchu
6 years ago
Theanswer has been derived below though you should know that kinetic energy of a particle performing SHM is maximum when x=0
Kinetic energy=\frac{kA^{2}}{2}-\frac{kx^{2}}{2} Maximum KE =\frac{kA^{2}}{2} =\frac{mw^{2}A^{2}}{2} =\frac{\rho Vw^{2}A^{2}}{2} =\frac{\rho V4 \pi ^{2}n^{2}A^ {2}}{2} (k=mw^{2} \boldsymbol{and} w. =2 \pi n)